OpenStudy (anonymous):

PLEASE HELP ME Retro Rides is a club for owners of vintage cars and motorcycles. Every year the club gets together for a ride. This year, 46 vehicles participated in the ride. The total number of tires of all the vehicles was 130. Assuming each car has 4 tires and each motorcycle has 2 tires, how many each of cars and motorcycles participated in the ride? 19 cars; 27 motorcycles 16 cars; 30 motorcycles 21 cars; 25 motorcycles 23 cars; 23 motorcycles

OpenStudy (danjs):

total tires = cars + motorcycles 130 = 4*C + 2*M 4 on each car, 2 on each motorcycle

OpenStudy (danjs):

Total vehicles = Cars + motorcycles 46 = C + H There are your two equations

OpenStudy (danjs):

130 = 4C + 2M 46 = C + M

OpenStudy (danjs):

19 cars , 27 motorcycles is what i got, would you like to see how?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

lol

OpenStudy (danjs):

ok you type this time... Solve the second equation for either C or M, it is up to you

OpenStudy (danjs):

46 = C + M, solve for C or M

OpenStudy (anonymous):

ty!